### v1

@@ 6.1.2 Experiments on $K$
Figure 9: RMSE of parameter estimates as a function of $K$, the number of parameter vectors drawn from the prior. Performance improves with larger $K$, but exhibits diminishing returns beyond $10{,}000$.
@@
Figure 9 shows how the RMSE of parameter estimates changes with increasing values of $K$. We observe that estimation error steadily decreases with larger $K$, but begins to plateau around $K=50{,}000$. Beyond this point, the marginal gains in accuracy are minimal, sometimes the error even seems to increase.
However, as shown in Figure 10, the training time increases rapidly with $K$. Based on these findings, we selected $K=10{,}000$ for our final model. This value offers a strong trade-off between training efficiency and estimation accuracy.
@@ 6.1.3 Experiments on neural network configuration
In parallel with the previous experiments, we investigated the impact of key architectural choices in the neural network design. While we did not extensively vary the number of hidden layers or neurons per layer, we fixed a standard architecture of three hidden layers with 32 neurons each for both the summary and inference networks.
Instead, we focused on the effect of three specific hyperparameters: the aggregation function, activation function, and loss function. This tuning phase was essential for understanding the estimator’s stability and sensitivity to design decisions.
For each configuration, we trained the model using $K=10{,}000$ parameter samples and $J=10$ datasets per parameter, as described earlier. After training, the network was evaluated on an independent test set of $1,000$ parameter samples drawn from the same prior.
@@
Agg. Function[RMSE] $\sigma^{2}$ $\alpha_{1}$ $\alpha_{2}$
@@
Act. Function[RMSE] $\sigma^{2}$ $\alpha_{1}$ $\alpha_{2}$
@@
Loss Function [RMSE] $\sigma^{2}$ $\alpha_{1}$ $\alpha_{2}$
MAE 0.064 6.23 6.42
Quantile loss 0.068 13.56 13.46
MSLE 0.062 5.81 5.89
TANH 1.19 13.54 13.77
Table 5: Root mean squared error (RMSE) for each parameter under different aggregation, activation, and loss functions. When testing one component, the others were held fixed.
@@
  • When testing loss functions, we used mean aggregation and Leaky ReLU activation.
In summary, based on the results of the simulation studies described in this appendix, we selected the following configuration for all experiments in the main text. We fixed the number of parameter draws to $K=10{,}000$ and the number of replicates per parameter to $J=10$, resulting in a training set of $100{,}000$ datasets. The input dimension $n_{t}$ was set to $1{,}440$, corresponding to one-minute returns over a full day. For the neural network architecture, we used three hidden layers with 32 neurons per layer in both the summary and inference networks. Finally, we adopted the mean aggregation function, Leaky ReLU as the activation function, and mean squared logarithmic error (MSLE) as the loss function. This configuration demonstrated strong performance across all models considered.

### v2

@@ 6.2.3 Experiments on $K$
Figure 11: RMSE of parameter estimates as a function of $K$, the number of parameter vectors drawn from the prior. Performance improves with larger $K$, but exhibits diminishing returns beyond 10,00010{,}000.
@@
Figure 11 shows how the RMSE of parameter estimates changes with increasing values of $K$. We observe that estimation error steadily decreases with larger $K$, but begins to plateau around K=50, [^0] K=50{,}000. Beyond this point, the marginal gains in accuracy are minimal, sometimes the error even seems to increase.
However, as shown in Figure 12, the training time increases rapidly with $K$. Based on these findings, we selected K=10, [^0] K=10{,}000 for our final model. This value offers a strong trade-off between training efficiency and estimation accuracy.
We also emphasize again that the choice of $K$ is ultimately model-dependent. In particular, the required $K$ tends to increase with the number of parameters being estimated and the overall complexity of the model. The simulation experiments reported here serve as benchmarks for specific Lévy processes, and may not directly generalize to all settings. For further guidance on selecting $K$, and additional simulation studies, we refer the reader to Sainsbury-Dale et al. [2023].
@@ 6.2.4 Experiments on Neural Network Configuration
In parallel with the previous experiments, we investigated the impact of key architectural choices in the neural network design.
We focused on the effect of two specific hyperparameters: the aggregation function and activation function. While we fixed a standard architecture of three hidden layers with 32 neurons each for both the summary and inference networks. This tuning phase was essential for understanding the estimator’s stability and sensitivity to design decisions.
For each configuration, we trained the model using K=10, [^0] K=10{,}000 parameter samples and $J=10$ datasets per parameter, as described earlier. After training, the network was evaluated on an independent test set of $1,000$ parameter samples drawn from the same prior.
@@
Agg. Function $\sigma^{2}$ $\alpha_{1}$ $\alpha_{2}$
@@
Act. Function $\sigma^{2}$ $\alpha_{1}$ $\alpha_{2}$
@@
Table 7: Root mean squared error (RMSE) for each parameter under different aggregation and activation functions. When testing one component, the others were held fixed.
@@
In summary, based on the results of the simulation studies described in this appendix, we selected the following configuration for all experiments in the main text. We fixed the number of parameter draws to K=10, [^0] K=10{,}000 and the number of replicates per parameter to $J=10$, resulting in a training set of 100,000100{,}000 datasets. The input dimension $n_{t}$ was set to 1,4401{,}440, corresponding to one-minute returns over a full day. For the neural network architecture, we used three hidden layers with 32 neurons per layer in both the summary and inference networks. Finally, we adopted the mean aggregation function, Leaky ReLU as the activation function, and mean squared logarithmic error (MSLE) as the loss function. This configuration demonstrated strong performance across all models considered.